OpenStudy (anonymous):

Debbie sells cookies and hotdogs. The daily cost of making cookies is $460 more than the difference between the square of the number of cookies sold and 20 times the number of cookies sold. The daily cost of making hotdogs is modeled by the following equation:

OpenStudy (anonymous):

C(x) = 2x2 - 80x + 1,200 C(x) is the cost in dollars of selling x hotdogs. Which statement best compares the minimum daily cost of making cookies and hotdogs? It is greater for hotdogs than cookies because the approximate minimum cost is $300 for cookies and $392 for hotdogs. It is greater for hotdogs than cookies because the approximate minimum cost is $360 for cookies and $400 for hotdogs. It is greater for cookies than hotdogs because the approximate minimum cost is $400 for cookies and $360 for hotdogs. It is greater for cookies than hotdogs because the approximate minimum cost is $392 for cookies and $300 for hotdogs.

OpenStudy (anonymous):

@liliegirl @Loser66 @lalaly

OpenStudy (anonymous):

@foreveryoung615 @mikel123

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (anonymous):

C(x) = 2x^2 - 80x + 1,200 This is the daily cost of hot dogs, where x is the number of hot dogs sold The daily cost of making cookies is $460 more than the difference between the square of the number of cookies sold and 20 times the number of cookies sold. The above sentence gives a similar function than C(x). Can you write that function? Lets say D(y)= ..................... where y is the number of cookies sold

OpenStudy (anonymous):

1200??

OpenStudy (anonymous):

The daily cost of making cookies is D(y) that is equal to $460 more than the difference between the square of the number of cookies sold and 20 times the number of cookies sold. Fairly complex sentence, can you translate English to maths?

OpenStudy (mathmale):

Baby: let x= the number of cookies sold. Then, how would you express "the square of the number of cookies sold'? How would you express "20 times the number of cookies sold"?

OpenStudy (anonymous):

d(y)=460^2*20 i dont really know

OpenStudy (mathmale):

Baby: your expression MUST include x, the number of cookies sold.

OpenStudy (anonymous):

You are getting somewhere, but it is still not the right way. Mathmale please use y instead of x. As x in this question already belongs to hot dogs. But mathmale has asked 2 very good questions to guide you. Then, how would you express "the square of the number of cookies sold'? How would you express "20 times the number of cookies sold"? y is the number of cookies sold

OpenStudy (anonymous):

20y?

OpenStudy (anonymous):

Yes, and the square of the number of cookies sold?

OpenStudy (anonymous):

the square for 460 is 21.4476106

OpenStudy (anonymous):

Well that is NO and NO. Do not mix the square of a number like three squared equals 3*3=9 or we write 3^2=9 or\[3^2=3*3=9\] once again do not mix this with the square root. \[\sqrt{9} =3\]

OpenStudy (anonymous):

i really dont understand this :/

OpenStudy (anonymous):

20 times the number of cookies sold You found out that this is 20*y or 20y For the square of the number of cookies sold That is gonna be y*y=y^2\[y^2\]

OpenStudy (anonymous):

I think what confuses you is this: The daily cost of making cookies is $460

OpenStudy (anonymous):

That sentence does not end there!!!!!

OpenStudy (anonymous):

The daily cost of making cookies is $460 MORE than the difference between the square of the number of cookies sold and 20 times the number of cookies sold.

OpenStudy (anonymous):

than the difference between the square of the number of cookies sold and 20 times the number of cookies sold. This can be translated into maths like this difference is - sign square of cookies sold is y^2 20 times the number of cookies sold is 20y All in one is y^2-20y Now the full sentence: The daily cost of making cookies is $460 MORE than the difference between the square of the number of cookies sold and 20 times the number of cookies sold. Daily cost we called D(y)=460 more than y^2-20y more than is + sign SO D(y)=460+y^2-20y

OpenStudy (anonymous):

The daily cost of hot dogs is given \[C(x) = 2x^2 - 80x + 1,200\] We found the daily cost of cookies \[D(y)=y^2-20y+460\]

OpenStudy (anonymous):

You can surely understand that the cost of any product has to depend on how many you sell.

OpenStudy (anonymous):

This was half of the question actually. Now we need to answer this: Which statement best compares the minimum daily cost of making cookies and hot dogs?

OpenStudy (anonymous):

i was thinking it was C

OpenStudy (anonymous):

This question asks about the minimal daily cost of hot dogs and cookies. This translated to maths: Find the minimal value of the above functions. Why C?

OpenStudy (anonymous):

sorry i have more than 1 question open i mean A because 300 looks to be the minimum

OpenStudy (anonymous):

baby_sb you are clearly guessing

OpenStudy (anonymous):

yes i am

OpenStudy (anonymous):

Well try to solve the problem first

OpenStudy (anonymous):

There are 2 ways to find the minimum value of a quadratic function. (quadratic means that the unknown is squared) One is to use algebra and make perfect squares. The second uses calculus and differentiation. Which one can you do?